Circular No. 8451

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2004 VR_{8} (LONEOS) An apparently asteroidal object discovered by LONEOS, designated

2004 VR_{8} (cf. MPS 118755, MPEC 2004-V48; discovery observation below), has been found to show a 10′′-diameter coma and a tail 16′′ long in p.a. 140^{o} on R-band CCD observations taken by C. W. Hergenrother (Lunar and Planetary Laboratory) with the 1.54-m Kuiper reflector at Catalina on Nov. 19.3 UT. Also, A. Nakamura (Kuma, Ehime, Japan) reports that 240-s unfiltered CCD frames taken with a 0.60-m reflector on Dec. 8.55 and 9.55 shows to object’s image to be slightly ‘softer’ than other field stars of similar brightness, and a possible very faint tail is visible to the southeast.

2004 UT  α_{2000}  δ_{2000}  Mag.
Nov. 3.34556  4^{h}06^{m}43.^{s}25  +13^{o}00′34′′.1  17.6

Additional astrometry, the following orbital elements, and an ephemeris appear on MPEC 2004-X36.

T = 2005 Sept. 2.6732 TT  ω = 63.^{o}1296
e = 0.510185  Ω = 71.2066 }2000.0
q = 2.375895 AU  i = 20.1326
a = 4.850593 AU n^{o} = 0.0922597 P = 10.68 years

COMETS C/2004 T6, T7, U4–U12 (SOHO) Additional SOHO Kreutz sungrazers (continuation to IAUC 8427; HO

= H. Otterstedt, XG = X. Gao):

Comet 2004 UT α_{2000} δ_{2000} Inst. F MPEC
C/2004 T6 Oct. 10.238 12^{h}47.^{m}0 − 8^{o}07′ C3/2 HS 2004-X23
C/2004 T7  13.431 12 49.4 − 9 59 C3/2 TH 2004-X23
C/2004 U4 17.288 13 22.8 −10 16 C2 BZ 2004-X23
C/2004 U5  21.096 13 30.2 −13 25 C3/2 HO 2004-X23
C/2004 U6  21.646 13 42.7 −12 20 C2  RK 2004-X23
C/2004 U7  22.413 13 41.8 −12 10 C2  XG 2004-X23
C/2004 U8  23.529 13 32.2 −13 59 C3/2 XL 2004-X24
C/2004 U9  24.929 13 36.8 −14 39 C3/2 TH 2004-X24
C/2004 U10  26.521 13 59.4 −13 47 C3/2 JS 2004-X24
C/2004 U11  26.904 13 42.6 −15 31 C2  HS 2004-X24
C/2004 U12  28.104 14 05.8 −14 20 C2  BZ 2004-X24
2004 December 10  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green